### v1

# Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging

Abstract
We uncover a new mechanism whereby the triple interplay of non-Hermitian pumping, bosonic interactions and nontrivial band topology leads to ultra-strong bosonic condensation. The extent of condensation goes beyond what is naively expected from the interaction-induced trapping of non-Hermitian pumped states, and is based on an emergent caging mechanism that can be further enhanced by topological boundary modes. Beyond our minimal model with 2 bosons, this caging remains applicable for generic many-boson systems subject to a broad range of density interactions and non-Hermitian hopping asymmetry. Our novel new mechanism for particle localization and condensation would inspire fundamental shifts in our comprehension of many-body non-Hermitian dynamics and opens new avenues for controlling and manipulating bosons.
Introduction.— In many-body physics, interactions between particles play a crucial role, driving interesting phenomena such as phase transitions [1, 2, 3, 4, 5], superconductivity [6, 7, 8], and quantum entanglement [9, 10, 11, 5]. Much of these physics can be observed through many-body dynamical simulations in both classical and quantum simulators [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
There has been growing excitement surrounding the interplay between many-body interactions and the non-Hermitian skin effect (NHSE) [29, 30, 31, 32, 33, 34, 35]. This excitement stems from the non-locality and sensitivity of the NHSE, which has a profound impact on particle localization [36, 37, 38, 39, 40, 41, 42] and condensation dynamics [43, 44, 45, 46, 42, 47], which can result in unexpected consequences when interplayed with quantum many-body interactions [29, 30, 31, 32, 33, 34, 35]. These include the emergence of a Fermi surface in a many-body non-Hermitian fermionic chain [29], exotic spin liquids from couplings to the environment [30], symmetry breaking and spectral structure of the interacting Hatano-Nelson model [31], and unconventional emission and absorption rates in non-Hermitian baths, as well as fundamental modifications of the quantum dynamics [32, 34]. Recently, the non-Hermitian Mott skin effect [48] has revealed the intriguing interplay between strong correlations and non-Hermitian topology in the many-body skin effect.
In this work, we uncovered a new mechanism whereby bosons can be made to condense very strongly through the triple interplay of non-Hermitian pumping, density interactions and non-trivial band topology. Due to the emergent non-locality of both the NHSE and the density interactions, the observed phenomenon is in stark contrast to the conventional behavior of interacting bosons, such as photons, which repel each other when they interact, leading to the quantum phenomenon known as photon blockade [49, 50, 51, 52]. Our simulations show that ultra-strong localization, and not just boundary state accumulation, is achieved when both non-Hermiticity and density interactions are present, which is not attainable by either alone.
Bosonic condensation from interactions and non-Hermiticity.– To demonstrate how non-Hermitian pumping, topology and bosonic interactions can interplay to cause unexpectedly strong bosonic condensation, we consider a minimal interacting 1D bosonic lattice model that contains topological edge modes at its boundary sites:
$H=$ $\sum_{x=1}^{L}t_{L}b_{2x-1}^{\dagger}b_{2x}+t_{R}b_{2x}^{\dagger}b_{2x-1}$ (1)
$+\sum_{x=1}^{L-1}\left(b_{2x}^{\dagger}b_{2x+1}+b_{2x+1}^{\dagger}b_{2x}\right)+\frac{U}{2}n_{x_{0}}^{2},$
which is a Su-Schrieffer-Heeger (SSH) lattice [53] with asymmetric intercell ($t_{L}$ and $t_{R}$) hoppings and unit intracell hoppings, equipped with a density interaction $\frac{1}{2}Un_{x_{0}}^{2}$ at site $x_{0}$, $U\geq 0$. Here $b_{x}$ ($b_{x}^{\dagger}$) is the bosonic annihilation (creation) operator at sites $x=1,2,\ldots,2L$, with $n_{x_{0}}=b_{x_{0}}^{\dagger}b_{x_{0}}$ the boson number operator. Non-Hermitian pumping [36, 54, 37, 55, 38, 56, 42, 57, 58, 59, 60, 44, 61, 40, 62, 63, 64, 65, 41, 66, 67, 68, 69, 70, 71, 72, 73, 47] occurs towards the left if $r=t_{L}/t_{R}>1$, accumulating the bosons towards the left boundary. However, when there are multiple indistinguishable bosons, this accumulation would also be affected by the repulsive ($U>0$) interaction at site $x_{0}$, with strength depending on bosonic occupancy. In this work, we have chosen the interaction $\frac{1}{2}Un_{x_{0}}^{2}$ that mimics the simplest possible nonlinearity in the mean-field limit, different from the usual Bose-Hubbard interaction [15, 49, 74, 50, 51, 52] by a local density shift [75, 76, 77, 78, 79].
Figure 1: Spatial density accumulation (a-d) and ultra-strong bosonic condensation (e) due to a boundary density interaction at $x_{0}=1$, for different two-boson initial states (1,1), (4,4), and (1,4) evolving via $H$ [Eq. 1]. $r=1$ or $4$ corresponds to Hermitian (a,b) or non-Hermitian (c,d) cases, and $U=0$ or $U\neq 0$ corresponds the absence (a,c) or presence (b,d) of the density interaction. (a-d) Heatmaps depict $\rho_{x}(t)$ [Eq. 3], the time-dependent boson number over the simulation duration $t\in[0,T]$ with $T=10$. The line plot above each heatmap shows the corresponding time-averaged boson number profile $\bar{\rho}_{i}$ [Eq. 4]. The bosons are most likely to be trapped at site $x=1$ in the presence of both leftwards NHSE ($r>0$) and nonzero interaction ($U>0$), as shown by the white and bright yellow regions in the heatmaps. While normally being considered as repulsive, the interaction term $U$ also acts as a trap at $x_{0}=1$ in non-equilibrium dynamics. (e) The dramatic interplay of NHSE pumping and bosonic interaction is most saliently expressed through the bosonic condensation at $x=1$, as given by the two-boson correlation $P_{11}$ [Eq. 5]. Here, we observe significantly larger $P_{11}$, indicated by the red curves in all three (e) subfigures, which demonstrates the greatly enhanced ability to trap both bosons at site 1 when both $U>0$ and $r>0$.
While one might naively expect a repulsive ($U>0$) density interaction to suppress the NHSE accumulation primarily, the observed behavior can be dramatically different even with just two bosons. Two distinct measures of particle accumulation can be defined: the spatial density accumulation and the many-body correlation. An initial 2-boson state $|\phi(t=0)\rangle$ evolves according to the Schrodinger equation $|\dot{\phi}(t)\rangle=-iH|\phi(t)\rangle$, and can be expressed as
$|\phi(t)\rangle=\frac{1}{\sqrt{2}}\sum_{x=1}^{2L}\sum_{x^{\prime}=1}^{2L}v_{xx^{\prime}}(t)|(x,x^{\prime})\rangle,$ (2)
where $v_{xx^{\prime}}(t)=v_{x^{\prime}x}(t)$ is the amplitude of the basis state $|(x,x^{\prime})\rangle=b_{x}^{\dagger}b_{x^{\prime}}^{\dagger}|0\rangle$ containing indistinguishable bosons at sites $x,x^{\prime}$. We define the 2-boson density at site $x=1,2,\ldots,2L$ by
$\rho_{x}(t)=\frac{\langle\phi(t)|b_{x}^{\dagger}b_{x}|\phi(t)\rangle}{\langle\phi(t)|\phi(t)\rangle},$ (3)
with $0\leq\rho_{x}(t)\leq 2$. Associated with it is the time averaged spatial density
[…]
where $T$ is the simulation duration. While $\rho_{x}(t)$ and $\bar{\rho}_{x}$ reveal where the bosons localize on the lattice, a high value of $\rho_{x}(t)$ or $\bar{\rho}_{x}$ can physically arise either due to strong single-boson localization at site $x$, or moderate double-boson condensation at the same site. To quantify this important distinction, we also examine the two-boson correlation probability
$P_{xx^{\prime}}(t)=\begin{cases}|v_{xx^{\prime}}(t)|^{2}&\text{if }x=x^{\prime},\\ |\sqrt{2}v_{xx^{\prime}}(t)|^{2}&\text{if }x\neq x^{\prime},\end{cases}$ (5)
which represents the probability of observing a boson at sites $x$ and $x^{\prime}$ at the same time $t$. For revealing underlying trends in the evolution of the two-boson correlation, we also define the time-smoothed correlation
[…]
which removes temporal oscillations shorter than a prescribed timescale $\Delta t$.
Figure 2: Extent of two-boson condensation due to a boundary density interaction at $x_{0}=1$ [Eq. 1], and its correspondence with the boundary localization of 2-boson spectral bands. (a) Time-averaged two-boson condensation probability $\bar{P}_{11}$ [Eq. 6] at site $x=1$, in the parameter space of non-Hermitian hopping asymmetry $r$ and density interaction strength $U$. (b) Corresponding two-boson spectra at $r=4$, which features five bands, with bands 2 and 4 being containing a pair of interaction-hybridized bulk and topological bosons. As $U$ increases, it creates a group of eigenstates at $E\approx U$ that exhibits strong localization $\psi_{11}$ (red) at $x=x_{0}=1$. Strong $\psi_{11}$ localization in the hybrid topological band 4 leads to suppressed $\bar{P}_{11}$ for the boundary-localized initial state $(1,1)$ (circled in dashed black), but enhanced $\bar{P}_{11}$ condensation for the initial state $(1,4)$ (solid black). By contrast, strong $\psi_{11}$ localization in the bulk bands 3 and 5 (circled in light blue) corresponds to enhanced $\bar{P}_{11}$ for the bulk initial state $(4,4)$.
We first consider scenarios where the density interaction is at the leftmost (boundary) site $x_{0}=1$. Figs. 1(a-d) showcase the evolution of the dynamical density $\rho_{x}(t)$ [Eq. 3] and its time-average $\bar{\rho}_{x}$ [Eq. 4] of three illustrative initial 2-boson state configurations. For the initial state $(1,1)$ (Top Row) where both bosons are already on $x_{0}=1$, the density interaction indeed repels the bosons in the Hermitian limit, as evidenced in the suppressed $\bar{\rho}_{x}$ density at $x=x_{0}=1$ across the Hermitian interacting vs. non-interacting cases (orange vs. blue). Somewhat expectedly, this suppression vanishes when the NHSE counteracts the repulsion by pumping the bosons towards $x_{0}=1$ (green and red). However, for the initial state $(4,4)$ (Center Row) where both bosons are initially far from $x_{0}=1$, the boson density oscillates and spreads out, failing to accumulate appreciably at $x_{0}=1$ except when both the non-Hermiticity and interaction are present (red). Indeed, in this non-equilibrium scenario, the $U>0$ density term behaves more like a local potential well that traps the $\rho_{x}(t)$ at $x=x_{0}$, rather than a repulsion. Qualitatively similar behavior is observed in the density evolution for initial state $(1,4)$ (Bottom Row), when one of the bosons is initially already at $x_{0}=1$ and is as such unaffected by the NHSE.
What is striking, however, is the ultra-strong two-boson condensation at site $x_{0}=1$ when both non-Hermiticity ($r>1$) and the density interaction ($U>0$) are present. This is revealed in the dynamical $P_{11}(t)$ [Eq. 5] plots of Fig. 1(e), which shows the probability of both bosons condensing at $x=1$. From the Top Row, when both bosons are prepared at $x=1$ (initial state =(1,1)), they can only remain both there if non-Hermiticity and interaction are simultaneously present (red). But most surprising is what happens when at least one boson is initially away from $x=1$ (Center and Bottom Rows): we observe overwhelmingly higher $P_{11}$ only when non-Hermiticity and interaction simultaneously interplay (red), compared to having either on their own (yellow, green). This suggests that the ability to trap both bosons and suppress photon blockade [50, 51, 49, 52] hinges on an emergent consequence of this interplay, a fact not evident from the density evolution alone [Fig. 1(a-d)].
To quantitatively characterize this two-boson condensation, we plot in Fig. 2(a) the time-averaged condensation probability $\bar{P}_{11}=\bar{P}_{11}(T)$ with $\Delta t=T$ [Eq. 6], which is the correlation $P_{11}$ smoothed over the entire simulation duration $T$. The heatmap of $\bar{P}_{11}$ in the interaction strength vs. non-Hermiticity $U$-$r$ plane [Fig. 2(a)] reveals significantly enhanced $\bar{P}_{11}$ (colored) for certain $(U,r)$ regions, compared to the $r=1$ Hermitian limit where $\bar{P}_{11}$ essentially vanishes (black). While $\bar{P}_{11}$ generally increases with $r$, optimal condensation $\bar{P}_{11}$ occurs at windows of $U$ that are highly dependent on the initial state, as will be explained below.
Figure 3: Caging mechanism of interaction term $U$ with initial state (2,2), left of interaction position $x_{0}=3$. (a) Snapshots of smoothed time-averaged two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) at $t=0,2.5,5,8$ for (a1) NH non-interacting and (a2) NH interacting cases. An inverted L-shape indicates an effective interaction range. (b1) Probability of both bosons at site 1. NH interacting case (red) shows optimal trapping. (b2) Single-boson probability ($\bar{P}_{1}$) at site 1 for comparison. Surprisingly, the interaction term acts as a barrier, confining bosons to sites 1-3 in the NH interacting case, demonstrating a unique caging effect not observed in single-boson scenarios or non-interacting cases. This highlights the distinctive role of interactions in two-boson dynamics. Lattice: $L=5$ cells; $U=1$; hopping asymmetry $r=4$; smoothing $\Delta t=2$.
Figure 4: Caging mechanism of interaction term $U$ with initial state (4,4), right of interaction position $x_{0}=3$. (a) Snapshots of smoothed time-averaged two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) at $t=0,1,1.6,2.2,3,8$ for (a1) NH non-interacting and (a2) NH interacting cases. Initial similar behavior in both cases ($t=0,1$). Interaction barrier effect emerges ($t=1.6,2.2$). NHSE-assisted confinement to left side in interacting case ($t=3,8$). (b1) Smoothed probability of both bosons at site 1 ($\bar{P}_{11}$ in Eq. 6). (b2) Single-boson probability ($\bar{P}_{1}$) for comparison. Notably, the NH interacting case demonstrates superior trapping ability for two bosons compared to the single-boson scenario, highlighting a unique interplay between interaction and non-Hermiticity. $U=6$ and smoothing $\Delta t=2$. (c) Time-averaged two-boson probability $\bar{P}{mn}$ (Eq. 6) within the cage, $\sum{m,n=1}^{x_{0}}\bar{P}_{mn}$, plotted against hopping asymmetry values $\log r\in[0,6]$ for various initial states: (c1) (4,4), (c2) (2,2), and (c3) (2,4). The orange curves ($U=1$) in (c1) and (c2) confirm the caging mechanism. A slight deviation for the orange curve ($U=1$) compared with the blue curve ($U=0$) is observed in (c3), attributed to insufficient interaction time between the two bosons. Lattice size $L=5$ unit cells.
Topological origin of ultra-strong bosonic condensation.– Intriguingly, for a density interaction at the $x_{0}=1$ boundary, it turns out that the ultra-strong boundary bosonic condensation is not just due to the interaction-NHSE interplay, but also relies crucially on SSH topology. This is evident from the 2-boson band structure plot in Fig. 2(b), where each eigenenergy $E$ is colored according to the overlap $\psi_{11}=\langle(1,1)|\psi\rangle$ of its corresponding eigenstate $|\psi\rangle$ with $|(1,1)\rangle=(b_{1}^{\dagger})^{2}|0\rangle$. Computed for our Hamiltonian $H$ [Eq. 1] at strong non-Hermitian hopping asymmetry $r=4$, it features 5 bands, with bands 2 and 4 arising from the hybridization of bulk and topological bands. This follows from the single-boson band structure [80], which contains two symmetrically gapped bulk bands separated by in-gap topological zero modes. The effect of the $U$ density interaction is to induce high $\psi_{11}$ overlap (red) successively from bands 3 to 5, as $U$ is increased.
To illustrate, with the initial state (1,1), suppressed $\bar{P}_{11}$at $U\approx 3$ corresponds to high $\psi_{11}$ overlap with band 4 [circled in black in Figs.2(a,b)]. This is because band 4 manifestly contain one bulk and one topological boson, which is not consistent with maintaining both bosons at $x_{0}=1$. As such, the same interaction strength of $U\approx 3$ leads to enhanced $\bar{P}_{11}$ (dashed black) for the initial state (1,4), where only one boson overlaps with the topological boundary mode. Likewise, $U\approx 1$ or $U\approx 5$ favors the $\bar{P}_{11}$ bosonic condensation for the bulk initial state (4,4), as circled in light blue, since the $\psi_{11}$ overlap with the bulk bands (3 and 5) are strongest at these interaction strengths $U$.
2-boson interactions as caging mechanisms.– Interestingly, the boson density interaction can also act as a non-local cage when the interaction position $x_{0}$ is not at the boundary site $1$. To demonstrate this, we shall show in Fig. 3 that, for an initial state with both bosons to the left of $x_{0}$, an interaction $U$ cages the evolved state, trapping it. And, more elaborately, we show in Fig. 4 that, for an initial state with both bosons to the right of $x_{0}$, the interaction will first act as a barrier to the evolved state, but will eventually still end up as a cage that traps the particle flux that leaked into it via NHSE pumping. We shall consider $x_{0}=3$ and initial states $(2,2)$ and $(4,4)$ in Fig. 3 and Fig. 4 respectively.
To illustrate the caging mechanism more intuitively, we express our 1D bosonic interacting model in its 2D configuration space $(x,x^{\prime})$, where $x$ and $x^{\prime}$ represent the positions of the two bosons. Here $(x,x^{\prime})$ and $(x^{\prime},x)$ refer to the same state due to bosonic statistics [74, 81, 82, 83, 77, 84]. A key observation is that the nonlinear term $\frac{1}{2}Un_{x_{0}}^{2}$ can be interpreted as a non-local potential offset [74, 77].
When the initial state is already inside the cage, bosons are predominantly confined within the cage (inverted-L shape) between sites $1$ and $3$, as shown in Fig. 4(a2). This contrasts with Fig. 4(a1), which shows no such suppression of spreading, despite the pumping direction being towards the left-top corner. When the initial state is outside the cage, as in Fig. 4(a), bosons are pumped to the left side of the lattice due to skin pumping. However, once they enter the cage, they become trapped. This is clearly demonstrated in Fig. 4(a2) at $t=3$ and $t=8$, where bosons are predominantly confined to the left side of the lattice, explaining the prominence of the red curve in Fig. 4(b1).
To prove that this caging mechanism is unique to the interacting scenario and not merely a consequence of an effective on-site potential, we conducted a comparative analysis. We examined the probability ($\bar{P}_{1}$, smoothed using the same method as in the two-boson case) of a single boson localized at the edge (site $1$) of a 1D SSH chain with an on-site potential $U$ at site $1$. The single-boson case is not as effective as the two-boson case, as shown in Fig. 3(b2) and Fig. 4(b2), which stand in stark contrast to the outstanding red curves in Fig. 3(b1) and Fig. 4(b1). This further proves that the interaction effect is unique to the two-boson case.
The efficacy of the caging mechanism is corroborated by analyzing the time-averaged two-boson probability within the caging region, $\sum_{m,n=1}^{x_{0}}\bar{P}_{mn}$, as a function of hopping asymmetry $\log r$ for diverse initial states (Fig.4(c)) in long-time dynamics over a period of 500. In Fig.4(c1) and (c2), the significantly higher values of the interacting case ($U=1$, orange curves) compared to the non-interacting case ($U=0$, blue curves) provide strong evidence for the caging phenomenon. Notably, Fig. 4(c3) exhibits a small deviation between the interacting and non-interacting scenarios, which we attribute to limited interaction time between the bosons when they start on opposite sides of the interaction site. This leads to a less pronounced caging effect, as the boson starting on site 2 is less likely to occupy site 3 due to the leftward skin pumping. This observation also explains the small difference between orange and blue curves in Fig. 3(b2) and Fig. 4(b2) – the caging effect is significant only when two bosons are interacting.
Conclusion.— To summarize, our investigation has revealed how the intriguing interplay between NHSE, topology and boson-boson interactions can unexpectedly lead to ultra-strong edge condensation of the bosons. The extent of this particle condensation goes beyond intuitive expectations of NHSE-pumped bosons trapped by the interaction, being also facilitated by the interaction-induced hybridization of topological and bulk states. Moreover, even more pronounced bosonic condensation can be induced by density interactions in the bulk ($x_{0}\neq 1$) due to an emergent caging mechanism. This trapping phenomenon is not observed in single-boson dynamics, and generalizes to higher numbers of bosons as well as generic density interactions that are non-local in the many-body configuration space. Our findings provide new perspectives on the role of interactions in the condensation and localization of bosons in non-Hermitian topological systems, offering a novel approach to trapping and controlling bosons. Our results are not limited to the specific model we have studied but can be extended to other systems with different forms of interactions and topologies. Experimentally, various platforms can potentially implement the setup for locally interacting bosons, including optical lattices [85, 86, 87, 18, 88, 89], photonic systems [90, 91, 92, 93, 94, 95, 75], and circuit QED [96, 25, 97, 98, 99]. Specifically, we have a potential experimental realization of the bosonic interacting model using resonator arrays in real space in the Supplemental Material [80].
Supplemental Material for
“Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging”
Mengjie Yang
Luqi Yuan
Ching Hua Lee
October 2, 2024
S1 Single-boson and two-boson spectra
In the main text, we have computed for our Hamiltonian $H$ [Eq. 1 in the main text] for two bosons at strong non-Hermitian hopping asymmetry $r=4$. The emergence of five distinct bands in the two-boson spectrum, compared to three bands typically seen in single-boson cases, requires careful explanation. Below, we elucidate the existence of these five bands by analyzing the system in the non-interacting limit ($U\to 0$).
The single-boson band structure, shown in Fig. S1(a), contains two symmetrically gapped bulk bands ($E_{1}$ and $E_{2}$) separated by in-gap topological zero modes ($E_{0}$). The corresponding eigenstates demonstrate distinct localization patterns, with extended states for bulk bands [Fig. S1(b)] and localized states for zero modes [Fig. S1(c)].
For the two-boson case in the non-interacting limit, the energy spectrum can be systematically constructed through all possible pairwise combinations of the single-particle energies. As shown in Fig. S2(a), this results in five distinct bands rather than three. The lowest [band 1 in Fig. S2] and highest [band 5 in Fig. S2] bands correspond to configurations where both bosons occupy the lower ($E_{1}\oplus E_{1}$) or upper ($E_{2}\oplus E_{2}$) bulk bands, respectively. The intermediate bands 2 and 4 in Fig. S2 arise from mixed occupations, where one boson occupies the zero mode while the other occupies either the lower ($E_{0}\oplus E_{1}$) or upper ($E_{0}\oplus E_{2}$) bulk band.
Figure S1: Single-boson spectrum of the SSH model. (a) The energy spectrum shows two bulk bands ($E_{1}$ and $E_{2}$) separated by topological zero modes ($E_{0}$). (b) Localization pattern $|\psi|^{2}$ of extended states in the bulk bands. (c) Localization of the topological zero mode, showing edge confinement.
Notably, the central band [band 3 in Fig. S2] emerges from the hybridization of two distinct configurations: one where the bosons are distributed between the lower and upper bulk bands ($E_{1}\oplus E_{2}$), and another where both bosons occupy the zero mode ($E_{0}\oplus E_{0}$). This hybridization explains why we observe five bands rather than six possible energy combinations, as the energetic proximity of these configurations leads to their mixing into a single hybridized band.
The localization patterns shown in Fig. S2(b) and (c) reflect this hybridization, exhibiting characteristics that can be understood as superpositions of the single-particle localization patterns.
S2 Possible experimental realization
In the main text, we have proposed a bosonic interacting model in a non-Hermitian SSH lattice, as shown in Eq. S1. The Hamiltonian is
$H=$ $\sum_{x=1}^{L}t_{L}b_{2x-1}^{\dagger}b_{2x}+t_{R}b_{2x}^{\dagger}b_{2x-1}$ (S1)
$+\sum_{x=1}^{L-1}\left(b_{2x}^{\dagger}b_{2x+1}+b_{2x+1}^{\dagger}b_{2x}\right)+\frac{U}{2}n_{x_{0}}^{2},$
Various physical platforms have experimentally realized non-Hermitian systems, including mature mechanical platforms [100, 101, 102, 103], photonic platforms [104, 45, 105, 106, 107], electrical circuit platforms [108, 109, 61, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120], and quantum circuit platforms [13, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130]. In the meantime, multiboson physics can be explored using cold atoms in optical lattices [15, 16, 17, 18, 19], photonic waveguide arrays [131, 132, 133, 134, 135, 136, 137, 138], and superconducting circuits [25, 26, 27, 28].
Below, we propose a potential experimental realization of the bosonic interacting model, as presented in Eq. S1, using resonator arrays in real space.
Arrays of micro-resonators have been extensively utilized to implement one-dimensional (1D) and two-dimensional (2D) photonic lattices, particularly in the investigation of non-Hermitian photonics over the past decade [139, 140, 141, 142, 143, 144]. These designs predominantly employ ring resonators, which can achieve exceptionally high Q factors [144].
It is noted that in Ref. [105], the authors demonstrate a feasible design of a one-dimensional non-Hermitian Su-Schrieffer-Heeger model based on photonic coupled resonant arrays in Fig. S3. The non-Hermitian asymmetric coupling can be realized by the judicious design of optical gain and loss elements into unidirectional coupling link rings, as shown in Fig. S3(b). The corresponding Hamiltonian is [105]
$H(k)=\left(\begin{array}[]{cc}0&t_{1}-\frac{\gamma}{2}+t_{2}e^{ik}\\ t_{1}+\frac{\gamma}{2}+t_{2}e^{-ik}&0\end{array}\right).$ (S2)
Here, in their model, the hopping amplitudes $t_{1}\pm\gamma$ and $t_{2}$ correspond to intracell hoppings $t_{L},t_{R}$ and intercell hopping (where we have set to be 1 for simplicity) in our model in Eq. S1, respectively. The non-Hermitian skin effect is observed in their model, which is consistent with our results.
Figure S2: Two-boson spectrum in the non-interacting limit ($U=0$). (a) Energy spectrum showing five bands arising from combinations of single-particle energies. Band 1 results from hybridization of $E_{1}\oplus E_{1}$. Band 2 results from hybridization of $E_{1}(E_{0})\oplus E_{0}(E_{1})$. Band 3 results from hybridization of $E_{1}\oplus E_{2}$ and $E_{0}\oplus E_{0}$ configurations. Band 4 results from hybridization of $E_{2}(E_{0})\oplus E_{0}(E_{2})$. Band 5 results from hybridization of $E_{2}\oplus E_{2}$. (b) Localization pattern of a zero-corner state in band 3. (c) Localization of mixed bulk-topological states in bands 2 and 4.
Figure S3: Schematic of the photonic coupled resonator arrays for realizing the non-Hermitian SSH model. The judicious design of optical gain and loss elements into unidirectional coupling link rings realise the non-Hermitian asymmetric coupling coupling coupling. This figure is adapted from Ref. [105].
Based on the design in Ref. [105], we only need the nonlinearity to be implemented in the system. The Kerr effect in the micro-resonators can introduce the nonlinearity. The Kerr effect, i.e., the third-order nonlinearity $\chi^{(3)}$, is a nonlinear optical effect that occurs when the refractive index of a material changes with the intensity of light.
The third-order nonlinearity can be achieved through natural materials (although the $\chi^{(3)}$ value is relatively small) or by employing cavity QED based on electromagnetically induced transparency (EIT) to achieve an equivalent $\chi^{(3)}$. The latter involves introducing atoms [79, 145, 146, 147] or quantum dots into a cavity [49, 148, 149, 150, 151], and through appropriate external field control, inducing a target optical field with strong effective nonlinearity. Additionally, the versatility of synthetic dimensions provides opportunities to implement the nonlinear term $\frac{U}{2}n_{x_{0}}^{2}$ in Eq. S1 demonstrated in Eq. (14) in Ref. [78].